# Mamba Neural Operator: Who Wins? Transformers vs. State-Space Models for PDEs

### v1

@@ 4 Experiments and Discussion
4.1 Dataset Description & Implementation Protocol
@@
PDEs Selection. We utilise datasets from PDEBench (Takamoto et al., 2022), a publicly available benchmark for partial differential equations (PDEs). We focus on three PDEs representing both stationary and time-dependent problems: Darcy Flow, Shallow Water 2D (SW2D), and Diffusion Reaction 2D (DR2D). All simulations are performed on a uniform grid. Detailed information about the datasets is provided in the Appendix.
Implementation & Evaluation Protocol. As Transformers have become the go-to architecture for PDE modelling and serve as the primary counterpart to SSM models, we selected three state-of-the-art Transformers as our baselines: GNOT (Hao et al., 2023), Galerkin Transformer (G.T.) (Cao, 2021), and OFormer (Li et al., 2022a). To achieve a fair comparison between Transformers and Mamba, we integrated the S6 block and Cross S6 block to replace self-attention and cross-attention in each model, creating modified versions of the original architectures. All three experimental methods initially adopt a linear attention mechanism as described in their original publications, while we evaluated two configurations for each of them: an implementation with standard softmax attention mechanism (w/S.A.) and a Mamba-enhanced implementation (our Mamba Neural Operator principle) (w/Mamba). All experiments were conducted on a single NVIDIA RTX 4090 GPU with 24GB of memory to ensure consistent and fair comparison conditions. Three metrics including Root Mean Squared Error (RMSE), Normalised RMSE (nRMSE), and Relative L2 Norm (RL2) were utilised for evaluation.
Table 1: Quantitative comparison on Darcy Flow ($\beta=1$) across three methods with linear attention (original version), softmax attention and Mamba. The performance is measured in terms of Root Mean Squared Error (RMSE), Normalised RMSE (nRMSE), and Relative L2 Norm (RL2), with the best-performing results highlighted.
DarcyFlow
[…]
GNOT (Hao et al., 2023) Galerkin 0.0070 0.0485 0.0370
w/S.A. Softmax 0.0061 0.0394 0.0299
w/Mamba (MNO) Mamba 0.0061 0.0367 0.0297
G.T. (Cao, 2021) Galerkin 0.0188 0.2027 0.1261
w/S.A. Softmax 0.0103 0.1050 0.0648
w/Mamba (MNO) Mamba 0.0061 0.0382 0.0286
OFormer (Li et al., 2022a) Normalised 0.0054 0.0253 0.0242
w/S.A. Softmax 0.0066 0.0324 0.0323
w/Mamba (MNO) Mamba 0.0054 0.0244 0.0241
Table 2: Quantitative comparisons on Shallow Water 2D (SW2D) and Diffusion Reaction 2D (DR2D) across three methods with linear attention (original version) and Mamba. The performance is measured in terms of Root Mean Squared Error (RMSE), Normalised RMSE (nRMSE), and Relative L2 Norm (RL2), with the best-performing results highlighted in green.
SW2D DR2D
[…]
GNOT (Hao et al., 2023)
Galerkin
0.0026 0.0025 0.0027 0.0567 0.6953 0.7233
w/Mamba (MNO) Mamba 0.0023 0.0022 0.0024 0.0060 0.0811 0.0570
G.T. (Cao, 2021)
Galerkin
0.0037 0.0035 0.0038 0.0083 0.1259 0.0723
w/Mamba (MNO) Mamba 0.0013 0.0013 0.0014 0.0012 0.0183 0.0099
OFormer (Li et al., 2022a) Normalised 0.0020 0.0020 0.0021 0.0177 0.2681 0.1559
w/Mamba (MNO) Mamba 0.0021 0.0021 0.0022 0.0123 0.1712 0.1134
4.2 Chose Your Winner: Transformer vs. Mamba for PDEs
We begin by evaluating the performance of Transformers, their variants, and Mamba on the Darcy Flow dataset, as presented in Table 1. The results demonstrate that incorporating Mamba consistently improves performance across all metrics and models. For GNOT, while the RMSE remains close, the nRMSE and RL2 values are reduced, indicating that Mamba effectively refines predictions. The G.T. sees the most significant enhancement, with the RMSE dropping by 40% when Mamba is used. This suggests that Mamba’s design addresses the shortcomings of traditional Galerkin-type attention in capturing complex PDE dynamics. For OFormer, Mamba not only retains the strong baseline performance but also achieves improvements across all metrics. The reduction in RL2 indicates that Mamba’s mechanism is better at mapping the solution space of PDEs with higher precision. Mamba also demonstrates an enhanced ability to capture the complex spatial correlations inherent to Darcy Flow more effectively.
The results across all datasets demonstrate a clear advantage of the Mamba Neural Operator overr architectures for PDEs. While Transformers are effective at capturing dependencies and patterns, Mamba’s specialised attention mechanisms provide a more understanding of the complex dynamics involved. By leveraging its unique cross-attention and self-attention blocks, Mamba not only achieves lower error rates but also enhances the stability and precision of predictions, particularly in highly nonlinear systems. These results suggest that Mamba enhances the expressive power and accuracy of neural operators, indicating that it is not just a complement to Transformers but a superior framework for PDE-related tasks, bridging the gap between efficient representation and accurate solution approximation.
We further validate Mamba’s potential through visualisations, as shown in Figure 2. The prediction and error maps reveal that Mamba consistently outperforms all Transformer variants, delivering more accurate solutions with lower error across challenging regions. Mamba handles fine details, particularly in capturing sharp gradients and subtle variations that standard attention mechanisms often miss. Compared to the Galerkin and Softmax attention Transformer models, Mamba reduces error propagation and improves spatial coherence. More Visualization result can be seen in Appendix C.
4.3 Why the Winner Wins: Breaking Down Mamba’s Win
Table 3: Comparisons with different
query positions using nRMSE.
Method Query Positions
Identical Diagonal
OFormer 0.0253 0.0318
w/S.A. 0.0324 0.0382
w/Mamba 0.0244 0.0314
Table 4: Comparisons with different dataset sizes using
nRMSE.
Method Dataset Sizes
9K 5K 2K 1K
GNOT 0.0485 0.0567 0.0777 0.1174
w/S.A. 0.0394 0.0400 0.0526 0.0776
w/Mamba 0.0367 0.0376 0.0481 0.0617
4.4 Ablation Study: Final Battles, Winner Takes All
Table 4 compares nRMSE performance across two query scenarios: Identical positions (input and query points are the same) and Diagonal positions (shifted inputs creating a mismatch). Experiments are done using Darcy Flow. While prior work shows that Transformers handle inconsistent input-query positions well (Li et al., 2022a), our results demonstrate a clear advantage of Mamba in both configurations. For Identical query positions, Mamba version achieves the lowest error, outperforming OFormer and its softmax variant, demonstrating Mamba’s superior ability to capture relationships when input and query points are perfectly aligned. For Diagonal query positions, where inputs and queries are misaligned, Mamba achieves the best performance compared to OFormer and its variant, demonstrating its superior ability to generalise under spatial shifts.
Table 4 compares nRMSE performance across different dataset sizes for GNOT, GNOT with softmax attention (w/ S.A.), and GNOT with Mamba. Experiments are carried out using Darcy Flow. As dataset size decreases, Mamba consistently achieves the lowest error, demonstrating superior performance and robustness in data-scarce scenarios. For instance, with the smallest dataset (1K), Mamba achieves an nRMSE of 0.0617, significantly lower than GNOT’s and GNOT w/ S.A.’s, showcasing its resilience and generalisation capability even with limited data. This highlights Mamba’s efficiency in learning meaningful representations with fewer data points, making it a powerful choice for real-world applications where data availability is a constraint.
5 Conclusion
Acknowledgments
CWC is supported by the Swiss National Science Foundation (SNSF) under grant number 20HW-1_220785. It also acknowledge CMI, University of Cambridge. JH was supported in part by the Imperial College Bioengineering Department PhD Scholarship and the UKRI Future Leaders Fellowship (MR/V023799/1). GY were supported in part by the ERC IMI (101005122), the H2020 (952172), the MRC (MC/ PC/21013), the Royal Society (IEC NSFC211235), the NVIDIA Academic Hardware Grant Program, the SABER project supported by Boehringer Ingelheim Ltd, Wellcome Leap Dynamic Resilience, and the UKRI Future Leaders Fellowship (MR/V023799/1). CBS acknowledges support from the Philip Leverhulme Prize, the Royal Society Wolfson Fellowship, the EPSRC advanced career fellowship EP/V029428/1, EPSRC grants EP/S026045/1 and EP/T003553/1, EP/N014588/1, EP/T017961/1, the Wellcome Innovator Awards 215733/Z/19/Z and 221633/Z/20/Z, CCMI and the Alan Turing Institute. AAR gratefully acknowledges funding from the Cambridge Centre for Data-Driven Discovery and Accelerate Programme for Scientific Discovery, made possible by a donation from Schmidt Futures, ESPRC Digital Core Capability Award, the EPSRC programme grant in ‘The Mathematics of Deep Learning’ EP/V026259/1, and CMIH and CCIMI, University of Cambridge.
Supplementary Material
In this supplementary material, we provide additional details regarding our methodology and a more comprehensive description of the dataset used in our experiments.
Appendix A Supplementary Information
A.1 Preliminaries: Transformer.
In each Transformer layer, an attention mechanism enables interaction between inputs at varying positions, followed by a position-wise fully connected network applied independently to each position. Specifically, the attention mechanism involves projecting an intermediate representation into three components—query $Q\in\mathbb{R}^{N\times d_{k}}$, key $K\in\mathbb{R}^{N\times d_{k}}$, and value $V\in\mathbb{R}^{N\times d_{v}}$—using three separate position-wise linear layers. These representations are then used to calculate the output as:
$\operatorname{Attention}(Q,K,V)=\mathrm{softmax}\bigg{(}\frac{QK^{T}}{\sqrt{d_{k}}}\bigg{)}V,$ (S.1)
of which the memory complexity is $O(n^{2})$. To reduce the computational inefficiency, Galerkin-type attention was proposed by [Cao, 2021] to remove Softmax attention with linear complexity. It defines as follows:
$\operatorname{Attention}_{g}(Q,K,V)=\frac{Q(\widetilde{K}^{T}\widetilde{V})}{d},$ (S.2)
in which $\widetilde{\cdot}$ denotes layer normalisation, as described in [Lei Ba et al., 2016]. The Galerkin-type attention mechanism involves two matrix product operations, resulting in a computational complexity of $O(nd^{2})$. This reduces the sequence length dependency to only $O(n)$.
A.2 Network Architecture.
As depicted in Figure 1 from the main paper, the data processing pipeline in our Mamba Neural Operator (MNO) is composed of three key stages: Bi-Directional Scan Expand, S6/Cross S6 Block, and Bi-Directional Scan Merge. When solving PDEs over a fixed grid, the input data can be structured as grid-based data, similar to an image. In the first stage, Bi-Directional Scan Expand, the MNO unfolds the input data into sequences by traversing the grid along two distinct paths. These sequences, representing input patches, are processed independently in the next step. The second stage, S6/Cross S6 Block, involves processing each patch sequence using either an S6 or Cross S6 block, depending on the model variation being employed. For instance, in the enhanced version of Mamba, the GNOT model utilises a Cross S6 block followed by an S6 block for further refinement. Finally, in the Bi-Directional Scan Merge stage, the processed sequences are reshaped and merged back together to generate the output map, completing the data forwarding process. This structured approach allows the MNO to efficiently handle grid-based input data, enabling scalable solutions for PDEs.
A.3 Definition of Cross S6.
Let $x$ and $x^{\prime}$ be two independent input vectors. Each input is processed through two independent linear transformation, resulting in corresponding parameter sets $(B,C,\Delta)$ for $x$ and $(B^{\prime},C^{\prime},\Delta^{\prime})$ for $x^{\prime}$. Specifically, these transformations are defined as:
$B,C,\Delta$ $=\text{Linear}_{x}(x),$ (S.3)
$\tilde{B}$ $=B+qB^{\prime},$ (S.4)
$\tilde{C}$ $=C+qC^{\prime},$
$\tilde{\Delta}$ $=\Delta+q\Delta^{\prime},$
Appendix B Further Implementation Details
@@
$\begin{cases}-\nabla\cdot\left(a(x,y)\nabla u(x,y)\right)=f(x,y),&\text{for }(x,y)\in\Omega,\\ u(x,y)=0,&\text{for }(x,y)\in\partial\Omega,\end{cases}$ (S.5)
@@
$S:a\mapsto u,\quad\text{for }(x,y)\in\Omega,$ (S.6)
with $a(x,y)$ and $u(x,y)$ as previously defined. Similar as PDEBench [Takamoto et al., 2022] protocol, we used only $\beta=1.0$ and we divided the training and testing ratio into 9:1 which contains 9,000 samples for training and 1,000 samples for testing.
@@
$\partial_{t}h+\partial_{x}(hu)+\partial_{y}(hv)$ $=0,$ (S.7)
@@
$h(t=0,x,y)=\begin{cases}2.0,&\text{if }\sqrt{x^{2}+y^{2}}<r,\\ 1.0,&\text{if }\sqrt{x^{2}+y^{2}}\geq r,\end{cases}$ (S.8)
@@
$S:h|_{t\in[0,t^{\prime}]}\mapsto h|_{t\in(t^{\prime},T]},\quad(x,y)\in\Omega,$ (S.9)
@@
Each sample in the dataset is discretized on a spatial grid of $128^{2}$ points and a temporal grid of 101 time steps. The first 10 time steps are used as input to the model, while the remaining 91 time steps serve as the target output. Following the protocol established by PDEBench [Takamoto et al., 2022], the dataset consists of 900 samples for training and 100 samples for testing.
@@
$\partial_{t}u$ $=D_{u}\partial_{xx}u+D_{u}\partial_{yy}u+R_{u},$ (S.10)
@@
$S:\{u,v\}_{t\in[0,t^{\prime}]}\mapsto\{u,v\}_{t\in(t^{\prime},T]},\quad(x,y)\in\Omega,$ (S.12)
@@
where $t^{\prime}=0.045\,\text{s}$ and $T=5.000\,\text{s}$, and the spatial domain is $\Omega=[-1,1]^{2}$. Here, $u=u(x,y,t)$ and $v=v(x,y,t)$ represent the activator and inhibitor, respectively. In this dataset, we follow the same discretization scheme similar to the Shallow Water equation, where each sample is downsampled to a spatial resolution of $128^{2}$ and a temporal resolution of 101 time steps (with 10 for input and rest of the 91 for target). Similar as the PDEBench protocol [Takamoto et al., 2022], the dataset includes 900 samples for training and 100 samples for testing.
Appendix C Extended Visual Results
This section extends the visual results presented in the main paper, providing a deeper comparison of the predictive performance and error distribution across different models. The supplementary figures illustrate the impact of incorporating Mamba on prediction accuracy and spatial coherence, further validating its advantages over traditional Transformer-based approaches.
Figure S.1 presents the prediction and error maps for the Galerkin Transformer (G.T.) and OFormer across three configurations: Galerkin attention, standard softmax attention, and Mamba (MNO). The results show that Mamba consistently achieves lower prediction errors, especially in regions with high variability, highlighting its ability to capture complex dynamics with greater precision compared to other configurations.
Figure S.2 and Figure S.3 provide visualised predictions over time for the Shallow Water and Diffusion Reaction datasets, respectively, using the original Galerkin Transformer and its Mamba-enhanced version. For the Shallow Water dataset, the Mamba-integrated model better preserves fine details and the circular wavefronts as time progresses, reflecting its superior capability to maintain spatial coherence. Similarly, in the Diffusion Reaction dataset, Mamba reduces the spread of error and better approximates the reference solution, demonstrating improved stability and generalisation in long-term simulations.
Overall, these visualisations clearly indicate that incorporating Mamba significantly enhances predictive accuracy and robustness, making it a superior choice for capturing intricate spatial-temporal patterns in complex PDE systems.
@@ Appendix C Extended Visual Results
Figure S.1: Results of prediction map and error map of the Galerkin Transformer and OFormer across three versions: Galerkin attention, Softmax attention, and Mamba.
@@
Figure S.2: Visualised prediction on Shallow Water dataset using Galerkin Transformer (G.T.) across the original and Mamba version.
@@
Figure S.3: Visualised prediction on Diffusion Reaction dataset using Galerkin Transformer (G.T.) across the original and Mamba version.

### v2

@@ IV Experiments and Discussion
IV-A Dataset Description & Implementation Protocol
@@
PDEs Selection. We utilise datasets from PDEBench[35], a publicly available benchmark for partial differential equations (PDEs). We focus on three PDEs representing both stationary and time-dependent problems: Darcy Flow, Shallow Water 2D (SW2D), and Diffusion Reaction 2D (DR2D). All simulations are performed on a uniform grid. Detailed information about the datasets is defined as follows:
@@
$\begin{cases}-\nabla\cdot\left(a(x,y)\nabla u(x,y)\right)=f(x,y),&\text{for }(x,y)\in\Omega,\\ u(x,y)=0,&\text{for }(x,y)\in\partial\Omega,\end{cases}$ (18)
@@
$S:a\mapsto u,\quad\text{for }(x,y)\in\Omega,$ (19)
with $a(x,y)$ and $u(x,y)$ as previously defined. Similar as PDEBench[35] protocol, we used only $\beta=1.0$ and we divided the training and testing ratio into 9:1 which contains 9,000 samples for training and 1,000 samples for testing.
@@
$\partial_{t}h+\partial_{x}(hu)+\partial_{y}(hv)$ $=0,$ (20)
@@
$h(t=0,x,y)=\begin{cases}2.0,&\text{if }\sqrt{x^{2}+y^{2}}<r,\\ 1.0,&\text{if }\sqrt{x^{2}+y^{2}}\geq r,\end{cases}$ (21)
@@
$S:h|_{t\in[0,t^{\prime}]}\mapsto h|_{t\in(t^{\prime},T]},\quad(x,y)\in\Omega,$ (22)
@@
Each sample in the dataset is discretized on a spatial grid of $128^{2}$ points and a temporal grid of 101 time steps. The first 10 time steps are used as input to the model, while the remaining 91 time steps serve as the target output. Following the protocol established by PDEBench[35], the dataset consists of 900 samples for training and 100 samples for testing.
@@
$\partial_{t}u$ $=D_{u}\partial_{xx}u+D_{u}\partial_{yy}u+R_{u},$ (23)
@@
$S:\{u,v\}_{t\in[0,t^{\prime}]}\mapsto\{u,v\}_{t\in(t^{\prime},T]},\quad(x,y)\in\Omega,$ (25)
@@
where $t^{\prime}=0.045\,\text{s}$ and $T=5.000\,\text{s}$, and the spatial domain is $\Omega=[-1,1]^{2}$. Here, $u=u(x,y,t)$ and $v=v(x,y,t)$ represent the activator and inhibitor, respectively. In this dataset, we follow the same discretization scheme similar to the Shallow Water equation, where each sample is downsampled to a spatial resolution of $128^{2}$ and a temporal resolution of 101 time steps (with 10 for input and rest of the 91 for target). Similar as the PDEBench protocol[35], the dataset includes 900 samples for training and 100 samples for testing.
Implementation & Evaluation Protocol. As Transformers have become the go-to architecture for PDE modelling and serve as the primary counterpart to SSM models, we selected three state-of-the-art Transformers as our baselines: GNOT[29], Galerkin Transformer (G.T.)[9], and OFormer[10]. To achieve a fair comparison between Transformers and Mamba, we integrated the S6 block and Cross S6 block to replace self-attention and cross-attention in each model, creating modified versions of the original architectures. All three experimental methods initially adopt a linear attention mechanism as described in their original publications, while we evaluated two configurations for each of them: an implementation with standard softmax attention mechanism (w/S.A.) and a Mamba-enhanced implementation (our Mamba Neural Operator principle) (w/Mamba). All experiments were conducted on a single NVIDIA RTX 4090 GPU with 24GB of memory to ensure consistent and fair comparison conditions. Three metrics including Root Mean Squared Error (RMSE), Normalised RMSE (nRMSE), and Relative L2 Norm (RL2) were utilised for evaluation.
TABLE I: Quantitative comparison on Darcy Flow ($\beta=1$) across three methods with linear attention (original version), softmax attention and Mamba. The performance is measured in terms of Root Mean Squared Error (RMSE), Normalised RMSE (nRMSE), and Relative L2 Norm (RL2), with the best-performing results highlighted.
DarcyFlow
[…]
GNOT[29] Galerkin 0.0070 0.0485 0.0370
w/S.A. Softmax 0.0061 0.0394 0.0299
w/Mamba (MNO) Mamba 0.0061 0.0367 0.0297
G.T.[9] Galerkin 0.0188 0.2027 0.1261
w/S.A. Softmax 0.0103 0.1050 0.0648
w/Mamba (MNO) Mamba 0.0061 0.0382 0.0286
OFormer[10] Normalised 0.0054 0.0253 0.0242
w/S.A. Softmax 0.0066 0.0324 0.0323
w/Mamba (MNO) Mamba 0.0054 0.0244 0.0241
TABLE II: Quantitative comparisons on Shallow Water 2D (SW2D) and Diffusion Reaction 2D (DR2D) across three methods with linear attention (original version) and Mamba. The performance is measured in terms of Root Mean Squared Error (RMSE), Normalised RMSE (nRMSE), and Relative L2 Norm (RL2), with the best-performing results highlighted in green.
SW2D DR2D
[…]
GNOT[29]
Galerkin
0.0026 0.0025 0.0027 0.0567 0.6953 0.7233
w/Mamba (MNO) Mamba 0.0023 0.0022 0.0024 0.0060 0.0811 0.0570
G.T.[9]
Galerkin
0.0037 0.0035 0.0038 0.0083 0.1259 0.0723
w/Mamba (MNO) Mamba 0.0013 0.0013 0.0014 0.0012 0.0183 0.0099
OFormer[10] Normalised 0.0020 0.0020 0.0021 0.0177 0.2681 0.1559
w/Mamba (MNO) Mamba 0.0021 0.0021 0.0022 0.0123 0.1712 0.1134
IV-B Chose Your Winner: Transformer vs. Mamba for PDEs
We begin by evaluating the performance of Transformers, their variants, and Mamba on the Darcy Flow dataset, as presented in Table I. The results demonstrate that incorporating Mamba consistently improves performance across all metrics and models. For GNOT, while the RMSE remains close, the nRMSE and RL2 values are reduced, indicating that Mamba effectively refines predictions. The G.T. sees the most significant enhancement, with the RMSE dropping by 40% when Mamba is used. This suggests that Mamba’s design addresses the shortcomings of traditional Galerkin-type attention in capturing complex PDE dynamics. For OFormer, Mamba not only retains the strong baseline performance but also achieves improvements across all metrics. The reduction in RL2 indicates that Mamba’s mechanism is better at mapping the solution space of PDEs with higher precision. Mamba also demonstrates an enhanced ability to capture the complex spatial correlations inherent to Darcy Flow more effectively.
@@
Figure 3: Results of prediction map and error map of the Galerkin Transformer and OFormer across three versions: Galerkin attention, Softmax attention, and Mamba.
@@
Figure 4: Visualised prediction on Shallow Water dataset using Galerkin Transformer (G.T.) across the original and Mamba version.
@@
Figure 5: Visualised prediction on Diffusion Reaction dataset using Galerkin Transformer (G.T.) across the original and Mamba version.
The results across all datasets demonstrate a clear advantage of the Mamba Neural Operator over Transformer architectures for PDEs. While Transformers are effective at capturing dependencies and patterns, Mamba’s specialised attention mechanisms provide a more understanding of the complex dynamics involved. By leveraging its unique cross-attention and self-attention blocks, Mamba not only achieves lower error rates but also enhances the stability and precision of predictions, particularly in highly nonlinear systems. These results suggest that Mamba enhances the expressive power and accuracy of neural operators, indicating that it is not just a complement to Transformers but a superior framework for PDE-related tasks, bridging the gap between efficient representation and accurate solution approximation.
We further validate Mamba’s potential through visualisations, as shown in Figure 2. The prediction and error maps reveal that Mamba consistently outperforms all Transformer variants, delivering more accurate solutions with lower error across challenging regions. Mamba handles fine details, particularly in capturing sharp gradients and subtle variations that standard attention mechanisms often miss. Compared to the Galerkin and Softmax attention Transformer models, Mamba reduces error propagation and improves spatial coherence.
Figure 3 presents the prediction and error maps for the Galerkin Transformer (G.T.) and OFormer across three configurations: Galerkin attention, standard softmax attention, and Mamba (MNO). The results show that Mamba consistently achieves lower prediction errors, especially in regions with high variability, highlighting its ability to capture complex dynamics with greater precision compared to other configurations.
Figure 4 and Figure 5 provide visualised predictions over time for the Shallow Water and Diffusion Reaction datasets, respectively, using the original Galerkin Transformer and its Mamba-enhanced version. For the Shallow Water dataset, the Mamba-integrated model better preserves fine details and the circular wavefronts as time progresses, reflecting its superior capability to maintain spatial coherence. Similarly, in the Diffusion Reaction dataset, Mamba reduces the spread of error and better approximates the reference solution, demonstrating improved stability and generalisation in long-term simulations.
IV-C Why the Winner Wins: Breaking Down Mamba’s Win
TABLE III: Comparisons with different query positions using nRMSE.
Method Identical Diagonal
OFormer 0.0253 0.0318
w/S.A. 0.0324 0.0382
w/Mamba 0.0244 0.0314
TABLE IV: Comparisons with different dataset sizes using nRMSE.
Method 9K 5K 2K 1K
GNOT 0.0485 0.0567 0.0777 0.1174
w/S.A. 0.0394 0.0400 0.0526 0.0776
w/Mamba 0.0367 0.0376 0.0481 0.0617
IV-D Ablation Study: Final Battles, Winner Takes All
Table III compares nRMSE performance across two query scenarios: Identical positions (input and query points are the same) and Diagonal positions (shifted inputs creating a mismatch). Experiments are done using Darcy Flow. While prior work shows that Transformers handle inconsistent input-query positions well[10], our results demonstrate a clear advantage of Mamba in both configurations. For Identical query positions, Mamba version achieves the lowest error, outperforming OFormer and its softmax variant, demonstrating Mamba’s superior ability to capture relationships when input and query points are perfectly aligned. For Diagonal query positions, where inputs and queries are misaligned, Mamba achieves the best performance compared to OFormer and its variant, demonstrating its superior ability to generalise under spatial shifts.
Table IV compares nRMSE performance across different dataset sizes for GNOT, GNOT with softmax attention (w/ S.A.), and GNOT with Mamba. Experiments are carried out using Darcy Flow. As dataset size decreases, Mamba consistently achieves the lowest error, demonstrating superior performance and robustness in data-scarce scenarios. For instance, with the smallest dataset (1K), Mamba achieves an nRMSE of 0.0617, significantly lower than GNOT’s and GNOT w/ S.A.’s, showcasing its resilience and generalisation capability even with limited data. This highlights Mamba’s efficiency in learning meaningful representations with fewer data points, making it a powerful choice for real-world applications where data availability is a constraint.
V Conclusion
Acknowledgments
CWC is supported by the Swiss National Science Foundation (SNSF) under grant number 20HW-1_220785. It also acknowledge CMI, University of Cambridge. JH was supported in part by the Imperial College Bioengineering Department PhD Scholarship and the UKRI Future Leaders Fellowship (MR/V023799/1). GY were supported in part by the ERC IMI (101005122), the H2020 (952172), the MRC (MC/ PC/21013), the Royal Society (IEC NSFC211235), the NVIDIA Academic Hardware Grant Program, the SABER project supported by Boehringer Ingelheim Ltd, Wellcome Leap Dynamic Resilience, and the UKRI Future Leaders Fellowship (MR/V023799/1). CBS acknowledges support from the Philip Leverhulme Prize, the Royal Society Wolfson Fellowship, the EPSRC advanced career fellowship EP/V029428/1, EPSRC grants EP/S026045/1 and EP/T003553/1, EP/N014588/1, EP/T017961/1, the Wellcome Innovator Awards 215733/Z/19/Z and 221633/Z/20/Z, CCMI and the Alan Turing Institute. AIAR gratefully acknowledges the support from Yau Mathematical Sciences Center, Tsinghua University.